OpenStudy (kydawg):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (perfecta):

sin (2x + 2x) sin2x cos 2x + cos2x sin2x sin2x cos2x + sin2x cos2x 2 sin 2x cos 2x sin 4x

OpenStudy (kydawg):

did you start on the left side?

OpenStudy (kydawg):

@PerfectA

OpenStudy (perfecta):

No Right

OpenStudy (kydawg):

once you get to sin4x idk what to do.

OpenStudy (welshfella):

Are you sure you've got the question right?

OpenStudy (welshfella):

i plugged in x = 15 degrees and got different values for each side

OpenStudy (kydawg):

What?

OpenStudy (dumbcow):

\[\cos(2x) = 1 - 2\sin^2(x)\] and \[\cos(4x) = 1-2\sin^2 (2x)\] Add them together and you get Right side

OpenStudy (welshfella):

AH yes. Sorry, I misread the question.